### v2

# From Source Properties to Strong-Field Tests: a multipronged analysis of GW250114 with an effective one-body model for generic orbits

Abstract
We present a detailed analysis of GW250114, the loudest gravitational-wave signal observed to date, using a waveform model capable of describing binary black holes in generic (eccentric and precessing) orbits. Our analysis builds on LIGO-Virgo-KAGRA (LVK)’s results, finding that the source is consistent at a probability of $\geq 96\%$ with the merger of two first-generation, nearly equal-mass, low-spin black holes, forming a remnant within the pair-instability mass gap. The signal’s high signal-to-noise ratio ($\gtrsim 75$) enables the detection of the subdominant $(4,\pm 4)$ multipoles, whose presence we confirm with higher evidence than previously reported by the LVK. Restricting the analysis even to post-peak data yields $\log_{10}\mathcal{B}\gtrsim 1$ in favor of models including the $(4,\pm 4)$ mode, demonstrating that this contribution remains detectable well into the post-merger phase. We further perform three independent tests of general relativity, complementary to those performed by the LVK: a modified residual analysis confirms that our semi-analytical model fully describes the signal without detectable discrepancies; a subdominant mode test finds that the amplitude of the $(4,\pm 4)$ multipoles agrees with general-relativistic expectations; and a parameterised analysis of the plunge-merger-ringdown regime recovers the GR expectation within the 50% credible region for the remnant mass and spin, and within the 90% interval for the $(2,\pm 2)$ peak amplitude. Collectively, these results reinforce GW250114 as a landmark event for a precision test of gravity.
Introduction: With an exceptional network signal-to-noise ratio (SNR) of $\gtrsim 75$, GW250114_082203 (GW250114, hereafter)—the loudest gravitational-wave (GW) signal of the first decade of GW astronomy—offers unprecedented statistical power for precision studies of binary black hole (BBH) mergers and for performing the most stringent GW-based tests of general relativity (GR) to date (Abac and others, 2025c, a; Kumar et al., 2025; Andrés-Carcasona and Caneva Santoro, 2025). Such a high-SNR signal exemplifies the remarkable sensitivity achieved by the ground-based GW detectors and the opportunities that future facilities promise (Cahillane and Mansell, 2022; Capote and others, 2025), particularly in the context of multi-phase tests of GR (Krishnendu and Ohme, 2021; Yunes et al., 2025). However, the statistical precision enabled by such signals places stringent demands on all aspects of the analysis. In particular, waveform accuracy becomes critical: even small modelling inaccuracies can bias the inferred source properties or masquerade as apparent deviations from GR (Q. Hu and J. Veitch (2023); R. S. Chandramouli, K. Prokup, E. Berti, and N. Yunes (2025); K. Chandra (2025); A. Gupta et al. (2024); A. Dhani, S. H. Völkel, A. Buonanno, H. Estelles, J. Gair, H. P. Pfeiffer, L. Pompili, and A. Toubiana (2025); A. G. Abac et al. (2025b); 47). Similarly, neglecting physical effects such as orbital eccentricity or spin-induced precession could in principle bias parameter estimates for certain sources.
The LIGO-Virgo-Kagra Collaboration (LVK) discovery and companion analyses established GW250114 as a quasi-circular, low-spin BBH merger with source-frame component masses $m_{1}=33.6^{+1.2}_{-0.8}\,M_{\odot}$ and $m_{2}=32.2^{+0.8}_{-1.3}\,M_{\odot}$, and spins $\chi_{1},\chi_{2}\leq 0.24,0.26$ (90% credible intervals) (Abac and others, 2025c, a). These estimates were shown to be robust across four state-of-the-art non-eccentric BBH waveform models (Varma et al., 2019; Pompili and others, 2023; Khalil et al., 2023; van de Meent et al., 2023; Ramos-Buades et al., 2023; Pratten and others, 2021; Colleoni et al., 2025). Using eccentric aligned-spin models, the LVK further constrained the orbital eccentricity to $e\leq 0.03$ at 13.33 Hz (Gamboa and others, 2025; Nagar et al., 2024), ascertaining the signal’s non-eccentric origin. Their post-inspiral analyses reported positive evidence for the $(4,\pm 4)$ subdominant mode ($\log_{10}\mathcal{B}=0.54^{+0.18}_{-0.18}$) and demonstrated that the post-peak emission is fully consistent with the Kerr $220$ quasi-normal mode (QNM) and its first overtone (Carullo et al., 2019; Isi et al., 2019). Residual analyses using best-fit waveforms and inspiral-merger-ringdown (IMR) consistency tests demonstrated agreement with GR, with the latter yielding constraints comparable to those obtained by hierarchically combining events in GWTC-4 (Abac and others, ) and proving that the event satisfied Hawking’s area increase law (Hawking, 1971). Finally, parameterised plunge-merger-ringdown tests within the pSEOBNR framework found both the $220$ and $440$ QNMs consistent with GR (Brito et al., 2018; Ghosh et al., 2021; Maggio et al., 2023; Pompili et al., 2025), while inspiral-phase tests based on post-Newtonian (PN) deviation parameters (Agathos et al., 2014) and principal-component analyses (Saleem et al., 2022) reinforced that GW250114 is a textbook example of a GR-consistent BBH signal.
Here we extend the LVK analyses by studying GW250114 with TEOBResumS-Dalí (Gamba et al., 2024; Nagar et al., 2025; Albanesi et al., 2025), a state-of-the-art Effective-One-Body (EOB) model capable of describing generic BBH orbits. We quantify the statistical evidence of the various physical effects modelled (eccentricity, spin precession or a combination of the two), demonstrate decisively the presence of the $(4,\pm 4)$ multipoles, and perform three independent tests of GR complementary to those presented by the LVK.
GR and beyond-GR models: A general relativistic BBH is intrinsically characterised by nine parameters: the mass ratio $q=m_{1}/m_{2}\geq 1$, the two (dimensionless) spins $\bm{\chi}_{1,2}$, eccentricty $e_{0}$ and mean anomaly $l_{0}$ at a reference frequency $f_{\rm ref}$. While the total mass does not affect the intrinsic dynamics, it determines the GW frequency and the signal’s loudness in physical units, making it relevant for detection and source property measurement.
TEOBResumS-Dalí semi-analytically models the radiation multipoles or GW modes $h_{\ell m}(t)$ of a BBH system through a two-stage approach within the EOB framework (Buonanno and Damour, 1999, 2000)  [^1] Our model version includes the $(\ell,m)=(2,\pm 2),(2,\pm 1),(3,\pm 3),(4,\pm 4)$ multipoles. The inspiral-plunge phase of the modes, $h_{\ell m}^{\mathrm{insp-plunge}}(t)$, is computed from the EOB dynamics via factorized and resummed PN expansions and numerical relativity (NR) calibrations, which capture the non-quasicircular effects during the final plunge (Damour et al., 2009; Damour and Nagar, 2007; Nagar et al., 2025). This is then smoothly connected to a phenomenologically constructed merger-ringdown template (Damour and Nagar, 2014; Del Pozzo and Nagar, 2017) $h_{\ell m}^{\mathrm{merger-RD}}(t)$ at mode-dependent matching times $t_{\ell m}^{\mathrm{match}}$ (Nagar et al., 2020):
$h_{\ell m}(t)=\begin{cases}h_{\ell m}^{\mathrm{insp-plunge}}(t)&t<t_{\ell m}^{\mathrm{match}}\\ h_{\ell m}^{\mathrm{merger-RD}}(t)&t\geq t_{\ell m}^{\mathrm{match}}~.\end{cases}$ (1)
where the latter combines inputs from black hole (BH) perturbation theory and NR simulations. TEOBResumS-Dalí differs from other EOB-based models (e.g. those presented in Refs. (Pompili and others, 2023; Khalil et al., 2023; van de Meent et al., 2023; Ramos-Buades et al., 2023; Gamboa et al., 2025; Gamboa and others, 2025)) in the choices made in its construction, from the resummations adopted to – most prominently – the treatment of non-quasicircular corrections and the merger-ringdown attachment procedure. For further details, we refer the reader to Gamboa and others (2025); Nagar et al. (2025)
To probe potential beyond-GR effects within the parameterised IMR framework, we employ two complementary approaches: (1) introduce fractional mode–amplitude deviation parameters, $\delta h_{\ell m}$, and constrain them directly from the data (Puecher et al., 2022; Gupta and others, 2025); (2) apply pTEOBResumS-Dalí (Chiaramello et al., 2025), which perturbs selected NR-calibrated parameters in TEOBResumS-Dalí’s dynamics and waveform generation. This second method is conceptually similar to the parameterised-deviation scheme in the pSEOBNR framework (Brito et al., 2018; Ghosh et al., 2021; Maggio et al., 2023; Pompili et al., 2025), used in LVK analyses of GW250114 and earlier events (Abbott and others, 2021, 2025) to constrain deviations in QNM frequencies and test the IMR-faithfulness of the ringdown within GR.
In our analysis, pTEOBResumS-Dalí focuses on three key parameters: $\delta M_{f}$ and $\delta\chi_{f}$, which modify the NR-informed remnant mass and spin, and $\delta A_{22}^{\mathrm{peak}}$, which alters the peak amplitude of the dominant $(2,\pm 2)$ mode. Changes in $\delta M_{f}$ and $\delta\chi_{f}$ affect the complex QNM frequencies of $h_{\ell m}^{\mathrm{merger-RD}}(t)$, leading to measurable shifts in the ringdown phase and damping times, while $\delta A_{22}^{\mathrm{peak}}$ perturbs the waveform amplitude near merger. Full details of the pTEOBResumS-Dalí implementation are given in Chiaramello et al. (2025).
Because TEOBResumS-Dalí models the inspiral–plunge with eccentric BBH dynamics refined by NR data, pTEOBResumS-Dalí can also quantify deviations from eccentric late-inspiral dynamics. However, TEOBResumS-Dalí currently transitions to a non-eccentric prescription for the merger–ringdown, assuming that binaries circularise sufficiently by this stage—a simplification we plan to address in future work. [^2] This limitation does not affect the results presented here, as we show in the following. As a result, pTEOBResumS-Dalí can measure deviations from a non-eccentric post-merger baseline.
Figure 1: Posterior distributions for orbital eccentricity $e$ and the waveform parameter $\zeta$ obtained with eccentric precessing (dash-dotted) and eccentric non-precessing (solid) models. Both analyses give consistent results, with median values $e\simeq 0.01$ and $\zeta\simeq 3.5$ for both cases.
Astrophysical nature of GW250114’s source: The LVK discovery and companion analyses established that GW250114 originated from a quasi-circular, low-spin BBH with negligible eccentricity (Abac and others, 2025c, a). Here we re-examine this astrophysical picture using an extended set of waveform hypotheses, to independently validate and quantify the LVK conclusions, and assess the robustness of the source characterisation across a broader modelling space. To do this, we analyse 8 s of publicly available Advanced LIGO data around GW250114 using four different BBH source hypothesis (see the Supplementary Material (SM) for details) (Abac and others, 2025d): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; (4) eccentric, precessing-spin.
Across all cases, the data favour a quasi-spherical, i.e., a non-eccentric but precessing configuration, consistent with the LVK findings, which didn’t report statistical evidence. Allowing for spin precession yields a slightly higher likelihood; removing precession reduces the maximum likelihood by $\Delta\max\ln\mathcal{L}\simeq-0.6$ and the Bayesian evidence by $\Delta\log_{10}\mathcal{B}=-0.12\pm 0.11$. Eccentric models—whether aligned-spin or precessing—perform worse, with $\Delta\max\ln\mathcal{L}\simeq-1$ and $\Delta\log_{10}\mathcal{B}\lesssim-1.55\pm 0.12$, and are thus more strongly disfavoured. Our eccentricity constraints also closely match those obtained by the LVK (Abac and others, 2025c, a). Figure 1 shows the joint posterior on orbital eccentricity $e$ and mean anomaly $\zeta$ obtained using uniform priors $e\in(0,0.4)$ and $\zeta\in(0,2\pi)$. We obtain $e=0.01^{+0.01}_{-0.01}$. Together with the Bayes factors quoted above, these results strengthen LVK’s claim that GW250114 is non-eccentric.
Figure 2: Posterior distributions of mass and spin parameters for GW250114. The 2D panels show the posteriors colored by the local posterior-to-prior density ratio $p(\bm{\xi}|d)/\pi(\bm{\xi})$, with warmer colors indicating greater support from the data. Black contours enclose 68% and 90% credible regions. The data favor an approximately equal-mass binary ($q\sim 1$) with low spin magnitudes ($\chi_{1,2}\sim 0.1$) and misaligned spins, as evidenced by the broad distributions of tilt angle cosines $\cos\theta_{1,2}$ measured relative to the orbital angular momentum.
Under the preferred (quasi-spherical) hypothesis, we recover source parameters consistent with the LVK’s estimates. We infer a source-frame total mass of $M_{T}=65.7^{+0.8}_{-0.9}\,M_{\odot}$ and mass ratio $q=1.03^{+0.06}_{-0.03}$, with low component spins, $\chi_{1}=0.13^{+0.11}_{-0.11}$ and $\chi_{2}=0.11^{+0.16}_{-0.09}$, as shown in Fig. 2. These correspond to individual masses of $m_{1}=33.3^{+0.9}_{-0.7}\,M_{\odot}$ and $m_{2}=32.2^{+0.7}_{-1.0}\,M_{\odot}$. Based on these estimates and data-driven GW-based population inference results (Abbott and others, 2023), we predict that the progenitors are likely first-generation (1G) BHs with $\geq 96\%$ certainty (see the SM for details).
The posteriors in $\cos\theta_{1,2}$ translate into spin-orbit tilt angles, defined with respect to the orbital angular momentum, of $\theta_{1}\in(105^{\circ},155^{\circ})$ and $\theta_{2}\in(92^{\circ},138^{\circ})$ at local posterior-prior density ratio $p(\bm{\xi}|d)/\pi(\bm{\xi})>15$, thereby corroborating the signal’s quasi-spherical BBH origin. These results are broadly consistent with LVK’s findings obtained using the NRSur7dq4 waveform model (see the SM for details).
From the posteriors and NR-informed fits (Hofmann et al., 2016; Jiménez-Forteza et al., 2017), we predict that the merger remnant has a final mass $M_{f}=63^{+1}_{-1}\,M_{\odot}\$and spin $\chi_{f}=0.67^{+0.01}_{-0.01}$. Using the surrogate fits of Varma et al. (2019, 2020), whose consistency with TEOBResumS-Dalí we have explicitly verified across the posterior-supported region (see the SM for details), we estimate that the remnant’s recoil speed is $v_{\mathrm{kick}}=141_{-117}^{+324}\,\mathrm{km/s}$. The remnant’s mass places it within the predicted pair-instability gap of $60-130\,M_{\odot}$ (Woosley and Heger, 2021; Farmer et al., 2019), and depending on the host environment, it may have either been ejected from or retained by it (Gerosa and Fishbach, 2021).
Figure 3: Posterior of azimuthal angle (measured in degrees at a reference frequency of 20 Hz) for various BBH models. The data favors a source with a bimodal azimuth separated by $\pi$, reflecting the binary’s symmetry under rotation $\phi\rightarrow\phi+\pi$.
Detectability and the temporal structure of the $(4,\pm 4)$ modes: The detectability of higher-order gravitational-wave multipoles depends on the source’s properties and orientation. For GW250114, the source’s near mass symmetry and negligible spins suppress the odd-$m$ modes, and the radiation is dominated by the $(2,\pm 2)$ component with no significant power mixing between modes of fixed $\ell$. Accordingly, based on the expected hierarchy of mode amplitudes, the $(\ell,m)=(4,\pm 4)$ mode is likely to be the strongest contributor after the quadrupole. This expectation is supported by the bimodal posterior in the azimuthal angle $\phi$, separated by $\pi$, reflecting the binary’s inherent $\phi\rightarrow\phi+\pi$ symmetry (see Fig. 3).
To ascertain the presence of the $(4,\pm 4)$ modes, we compare the full–multipolar TEOBResumS-Dalí model with an otherwise identical version excluding these harmonics. The run including $(4,\pm 4)$ yields a higher maximum log-likelihood, with $\Delta\ln\mathcal{L}_{\max}\simeq 11.5$, and a modest Bayes factor $\log_{10}\mathcal{B}\simeq 0.66$, with the maximum SNR gain ($\approx 4.79$) being comparable to the “optimal” SNR upper limit obtained for these modes in Abac and others (2025c) using the formalism of Mills and Fairhurst (2021). The modest Bayes factor, despite the substantial log-likelihood increase, reflects that the higher-harmonic improvement is restricted to a narrow region of parameter space and is therefore penalised by model complexity.
Figure 4: Time-gated analysis of GW250114 to test the contribution of the $(4,\pm 4)$ mode as a function of the start time $\Delta t^{\mathrm{geo}}_{\mathrm{start}}/M$, measured relative to the signal peak in units of remnant black-hole mass. The curves show the change in the log Bayes factor $\Delta\log_{10}\mathcal{B}$ (colored) and the improvement in matched-filter SNR relative to a quadrupole-only model (black). The $(4,\pm 4)$ contribution remains detectable even post signal’s peak, thereby showing that GW250114 is the first GW event with a measurable $(4,\pm 4)$ contribution persisting into the post-merger phase.
We strengthen this finding by identifying when this multipole makes the most significant contribution to the signal. To this end, we use a time-domain likelihood (see SM), progressively excluding data before a start time $\Delta t^{\mathrm{geo}}_{\mathrm{start}}$ (Chandra and Calderón Bustillo, 2025), defined relative to the signal peak in units of the remnant black-hole mass. At each time step, we compare the TEOBResumS-Dalí model in its quasi-spherical configuration, including both the quadrupole and $(4,\pm 4)$ multipole, against its quadrupole-only analogue. We fix the sky location and peak time $t_{\mathrm{peak}}^{\mathrm{geo}}$ to their maximum-likelihood values for this analysis. As shown in Fig. 4, the $(4,\pm 4)$ multipole has a detectable imprint even in the post-peak portion of the signal, up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}/M=5$, beyond which the maximum matched-filter SNR relative to a quadrupole-only model drops below 3.
Moreover, as the plot shows, restricting to late-time data yields $\log_{10}\mathcal{B}\sim 1-4$ against the quadrupole-only hypothesis, larger than the one quoted from the full inspiral-merger-ringdown analysis. This is because in a full signal analysis, the earlier inspiral dilutes the relative information carried by the $(4,\pm 4)$ mode. These values are also larger than those obtained from the KerrPostmerger analysis of Abac and others (2025c); we attribute this difference primarily to prior choices and model features.  [^3] KerrPostmerger samples over each mode’s phase at merger independently, whereas in our framework these phases are fixed by the inspiral–plunge dynamics (Gennari et al., 2024). Additionally, KerrPostmerger employs a broader mass prior.444pSEOBNR analyses indirectly suggest such a contribution by constraining deviations in the 440 QNM frequencies, and by comparing the results with simulated signals in which the $(4,\pm 4)$ multipole is removed, although—as quoted in the paper—“the analysis cannot by itself establish whether the 440 QNM is present in the data”.
Figure 5: Residual distributions for Hanford (left) and Livingston (right) detectors after subtracting the whitened TEOBResumS-Dalí posterior waveforms from whitened strain data. Individual curves show posterior samples, colored by their Jensen–Shannon divergence from the expected white noise distribution (black dashed). We find that the residuals are statistically indistinguishable from noise ($\text{JSD}=0.008\pm 3\times 10^{-4}$, $p>0.05$ for all samples), thereby indicating that the data is well described by GR-consistent BBH waveforms.
Figure 6: Posterior for the fractional deviations in the final mass, $\delta M_{f}$, and final spin, $\delta\chi_{f}$, relative to the GR predictions for the quasi-spherical, quasi-circular, and eccentric, non-precessing model. The outer contours enclose the $90\%$ credible regions and in both non-eccentric cases the GR expectation $\delta M_{f}=\delta\chi_{f}=0$ lies within 50% credible level. For comparison, we also show the corresponding constraints derived from the publicly available pSEOBNR (LVK) samples (dashed dotted blue). To enable a direct comparison, we first map the pSEOBNR BBH posteriors to obtain the IMR-predicted quantities, and separately map the deviated fundamental mode QNM spectra samples to their corresponding remnant properties. From these, we compute the fractional deviations. As shown, the pSEOBNR posteriors also include the GR value at their $50\%$ credible interval. Our constraints are tighter, reflecting the additional information contributed by the subdominant angular mode present in our analysis.
Signal’s consistency with GR: We begin by assessing the consistency of the data with the GR prediction using a modified residual analysis. The LVK companion paper performed this test by subtracting the maximum-likelihood waveform from the data and showing that the residual network SNR is compatible with the noise background with a Kolmogorov–Smirnov p-value of 0.34. Here we complement that study by performing a slightly different test: instead of subtracting a single waveform, we subtract the full ensemble of posterior-supported quasi-spherical TEOBResumS-Dalí waveforms from the data. This approach avoids relying on the single maximum-likelihood waveform returned by an off-the-shelf stochastic sampler, which is not optimised to locate the true maximum log-likelihood point in high-dimensional parameter space (Speagle, 2020).
Figure 5 shows the resulting whitened residuals, obtained by subtracting the whitened posterior waveforms from the whitened strain and comparing the distribution with a zero-mean univariate Gaussian. Across the posterior ensemble, the residuals are statistically indistinguishable from the expected coloured Gaussian noise. We measure Jensen–Shannon divergences of $0.088\pm 0.003$ and Kolmogorov-Smirnov $p$-values of $0.23\pm 0.24$, all satisfying $\mathrm{JSD}<8\times 10^{-3}$ and $p>0.05$. These show that the general-relativistic TEOBResumS-Dalí waveform model fully captures the gravitational-wave signal and the remaining fluctuations are consistent with detector noise.
Figure 7: Posterior distributions of the fractional $(4,4)$-mode deviation, $\delta h=(1+\delta A_{44}e^{i\delta\phi_{44}})$. The blue curve shows the observed event, the pink curve a NRSur7dq4 synthetic signal recovered with TEOBResumS-Dalí in zero noise, and the green curve a matched $TEOBResumS-Dalí$ recovery. In all cases, the $(4,4)$ mode deviation is consistent with GR, yielding a posterior peaked near the value of 11 (vertical black line). However, in the event and the mismatched synthetic signal, we observe a slight shift of the posterior amplitude to the right, likely due to waveform systematics.
Assuming that the dominant quadrupole $(\ell,m)=(2,\pm 2)$ mode agrees with GR, we next test the subdominant $(4,\pm 4)$ multipole under a quasi-circular signal hypothesis.  [^5] As shown earlier, precessing and non-precessing configurations are statistically indistinguishable for this event, supporting the use of the simpler non-precessing analysis. Also, precession leads to significant power mixing between modes of fixed $\ell$, indicating one needs to add these deviations in the co-precessing frame, which we don’t. Following and extending Puecher et al. (2022); Gupta and others (2025), we introduce a time-independent fractional mode deviation, parameterised by an amplitude and a phase deviation via $h_{44}\ \rightarrow\ h_{44}\,(1+\delta A_{44}e^{i\delta\phi_{44}})$, and constrain it directly from the data while enforcing $h_{\ell,-m}=(-1)^{\ell}h_{\ell m}^{\ast}$. Figure 7 shows the deviations posteriors, which are consistent with the GR prediction of $(1+\delta A_{44}e^{i\delta\phi_{44}})=1$ at the 90% credible level, though the amplitude posterior shows a long tail towards higher values and a shift of the peak away from the GR expectation. To test the robustness of this finding, we perform two additional zero-noise simulations with a GW250114-like signal: (i) matched $TEOBResumS-Dalí$ recovery, and (ii) unmatched NRSur7dq4TEOBResumS-Dalí recovery. The former yields a unimodal posterior peaked at the GR value, while the latter reproduces the shift in in peak amplitude deviation similar to that observed in the event analysis. This suggests that waveform systematics are responsible for this feature. To our knowledge, this is one of the first attempts to directly test the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR in a BBH signal.
Next, we perform two complementary parameterised tests of the plunge–merger–ringdown regime. Our first test introduces fractional deviations in the remnant mass and spin, $(\delta M_{f},\delta\chi_{f})$, relative to their GR values, sampling both from uniform priors over (0.2,0.2)(-0.2,0.2). Figure 6 shows the resulting posteriors for three non-precessing signal hypotheses: quasi-spherical (dash-dotted rose), quasi-circular (dashed teal), and eccentric (solid black). In both non-eccentric cases, the GR point $(\delta M_{f}=\delta\chi_{f}=0)$ lies well within the $\sim 50\%$ credible region while for the eccentric case it lies within the $70\%$ credible region. The corresponding Savage–Dickey ratios, $\log_{10}[p(\delta M_{f},\delta\chi_{f}|d)/\pi(\delta M_{f},\delta\chi_{f})]=1.51\mathrm{~and~}2.84$, provide positive support for GR-consistent post-inspiral dynamics for both non-eccentric signal source hypothesis.
Additionally, we evaluate the Bayesian support for GR-consistent plunge–merger–ringdown against beyond-GR extensions. For quasi-circular binaries, we obtain $\Delta\log_{10}\mathcal{B}^{\rm GR}_{\rm bGR}=2.30\pm 0.12$ ($\mathcal{B}\simeq 200$), constituting decisive evidence for GR. Although the maximum likelihood improves only marginally ($\Delta\ln\mathcal{L}_{\rm max}=0.49$), the deviation models are penalised by their larger prior volume. Quasi-spherical binaries show similar behaviour, with $\Delta\log_{10}\mathcal{B}=2.63\pm 0.12$ and $\Delta\ln\mathcal{L}_{\rm max}=0.57$, again strongly favouring GR. Taken together, both the metrics: the Savage-Dickey ratios and the Bayes factors indicate no evidence for a GR-inconsistent BBH plunge–merger–ringdown.
As previously mentioned, this analysis, though similar in spirit to the pSEOBNR test performed by the LVK (Brito et al., 2018; Pompili et al., 2025; Abac and others, 2025a), differs in its implementation. pSEOBNR constrains deviations from the IMR-expected QNM frequencies, while we directly constrain deviations in the IMR-expected remnant properties that inform the QNM frequencies that is used for modelling the post-peak waveform. Therefore, to compare the results, we use the LVK’s publicly available pSEOBNR posterior samples and reconstruct the corresponding remnant properties. Starting from the intrinsic binary parameters, we use NR fitting formulae to obtain the GR-consistent inspiral predictions for the remnant, $\bigl(M_{f}^{\rm IMR}(1+z),\,\chi_{f}^{\rm IMR}\bigr)$. Independently, we map the measured quasi-normal–mode quantities $\bigl(f_{220}(1+\delta f_{220}),\,\tau_{220}(1+\delta\tau_{220})\bigr)$ to the corresponding remnant mass and spin, $\bigl(M_{f}^{\rm dev}(1+z),\,\chi_{f}^{\rm dev}\bigr)$, using the relations of Berti and Cardoso (2006). We then quantify departures from GR through $\delta X_{f}\equiv X_{f}^{\rm dev}/X_{f}^{\rm IMR}-1,$ for $X_{f}\in\{M_{f},\chi_{f}\}$. As shown in Fig. 6, the resulting pSEOBNR constraints include the GR prediction also within the $50\%$ credible interval. Our bounds are somewhat tighter, consistent with the fact that our analysis leverages information from the subdominant $(4,\pm 4)$ multipole in addition to the dominant mode.
Figure 8: Posterior distribution of the fractional deviation in the peak amplitude of the dominant $(2,2)$ mode, $\delta A_{22}^{\rm peak}$. The outer dashed blue contour encloses the 90% credible region, while the dashed black line marks the median. The GR expectation, $\delta A_{22}^{\rm peak}=0$, lies at the 82% credible level, indicating consistency with GR.
Our second test allows a fractional deviation in the peak amplitude of the dominant $(2,2)$ mode, $\delta A_{22}^{\rm peak}$, sampling from a uniform prior on (1,4)(-1,4) under the quasi-circular assumption. For such an assumption, symmetry imposes the relation $h_{\ell,-m}=(-1)^{\ell}h_{\ell,m}^{\ast}$, ensuring that the (2,2)(2,-2) mode is fully determined by the $(2,2)$ mode and removing additional freedom in the waveform morphology. We find that $\delta A_{22}^{\rm peak}=0$ lies within the $\sim$82% credible interval, consistent with GR (See Fig. 8) and has a posterior width of $\mathcal{O}(10^{-1})$. For comparison, this value is just one to two orders of magnitude larger than the typical systematic uncertainty in NR simulations obtained from different codes (see e.g. Fig. 8 of Rashti et al. (2025)). This fact is rather remarkable, showcasing how GW observations are soon to enter the high precision regime.
Discussion: Our complementary analysis confirms and strengthens Abac and others (2025c, a) results, indicating GW250114 as a nearly mass-symmetric quasi-spherical BBH merger with total mass $M_{T}\simeq 66\,M_{\odot}$ and low spins $\chi_{1,2}\simeq 0.1$, decisively disfavouring eccentric configurations ($\mathcal{B}\sim 20$). These properties point to a first-generation binary formed through isolated field evolution (Postnov and Yungelson, 2014) with $\geq 96\%$ probability. Additionally owing to these properties and the signal’s high SNR ($\gtrsim 75$), we obtain decisive evidence for the subdominant $(4,\pm 4)$ multipoles, detectable up to $\Delta t^{\mathrm{geo}}_{\mathrm{start}}=5M$ beyond the signal’s peak, thereby strengthening the finding of Abac and others (2025a).
We then subject GR to a trio of stringent, complementary tests in the strong-field, non-linear regime with this signal. First, we perform a residual analysis, subtracting the full ensemble of posterior-supported waveforms from the data, and confirm that TEOBResumS-Dalí fully captures the signal, with no statistically significant deviations from noise ($p\simeq 0.3$). Second, we carry out the one of the first direct parameterised test of the subdominant $(4,\pm 4)$ multipole’s amplitude consistency with GR  in a BBH signal, introducing a fractional mode deviation and recovering its posterior from the data. We find that the GR value lies within the 90% credible region, but the deviation posterior exhibits a mild shift in amplitude, a structure reproduced in simulated signal analysis where a NRSur7dq4 waveform is recovered using TEOBResumS-Dalí, but not in matched $TEOBResumS-Dalí$ recoveries. The agreement between the signal and mismatched simulated signals points to waveform systematics in the $(4,\pm 4)$ mode of TEOBResumS-Dalí. Lastly, using the first parameterised eccentric BBH waveform model, we find a GR-consistent plunge–merger–ringdown portion, with the GR expectation $(\delta M_{f}=\delta\chi_{f}=0)$ lying well within the 50% credible region for both non-eccentric cases and $\delta A^{\mathrm{peak}}_{22}$ lying inside the 90% credible region. Regarding the latter, we note that while finalising the paper, we became aware of Grimaldi and others (2025) that also constrains the GW250114 merger amplitude.
Collectively these analyses, together with those in Abac and others (2025c, a), provide compelling and multifaceted evidence for the validity of GR in the dynamical, strong-field gravity of a BBH coalescence. GW250114 thus stands as a watershed event, heralding a new era of precision strong-gravity physics where future high-SNR detections will enable even more stringent constraints on potential deviations from GR.
Acknowledgements: The authors thank Arnab Dhani, Ish Gupta, B. Sathyaprakash, and Juan Calderón Bustillo for their comments and valuable suggestions, which greatly improved this paper. They also appreciate Alessandra Buoananno’s and Elisa Maggio’s comments and constructive feedback. RG also thanks K. Glenn for help in cleaning the manuscript, while KC thanks A. Hussain. KC acknowledges support from NSF grants AST-2307147, PHY-2207638, and PHY-2308886. RG acknowledges support from NSF Grant PHY-2020275 (Network for Neutrinos, Nuclear Astrophysics, and Symmetries (N3AS)) This work has made use of data obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Finally, the authors acknowledge using the Nemo cluster at the University of Wisconsin-Milwaukee (supported by NSF grants PHY-1626190 and PHY-2110594) and the Gwave cluster at Pennsylvania State University (supported by NSF grants OAC-2346596, OAC-2201445, OAC-2103662, OAC-2018299, and PHY-2110594) for the computational analyses presented here.
Supplementary Material
Analysis Setup:
We analyse 8 s of publicly available Advanced LIGO data around the event (Abac and others, 2025d), sampled at 4096 Hz, using Bilby with the dynesty sampler and the TEOBResumS-Dalí waveform model in four different configurations (Ashton and others, 2019; Speagle, 2020): (1) non-eccentric, aligned-spin; (2) non-eccentric, precessing-spin; (3) eccentric, aligned-spin; and (4) eccentric, precessing-spin BBH mergers. This model-selection step identifies the most probable configuration a priori, thereby maximising the constraining power of subsequent deviation measurements  [^6] By picking the most probable configuration, we shrink the allowed parameter space, which makes the subsequent measurements give tighter constraints..
Following Romero-Shaw and others (2020), we adopt isotropic priors for sky position and orientation, and a distance prior uniform in comoving volume and source time. Configuration-specific priors are used for intrinsic parameters–particularly spins, eccentricity, and mean anomaly–evaluated at a reference frequency of 13.3 Hz. To probe deviations from GR, we use settings similar to those above, introducing uniform priors on the relevant deviation parameters as needed.
For our full-signal analyses, we use the Whittle likelihood, with the publicly available data power spectrum estimated on-source via a Bayesian approach that combines splines for broadband features and Lorentzians for narrow spectral lines (Cornish and Littenberg, 2015). For measuring the persistence of the subdominant $(4,\pm 4)$ multipole in the post-inspiral signal, we instead employ a time-domain likelihood implemented in Bilby (Ashton and others, 2019); further details are provided below.
The likelihood functions:
Bayesian inference libraries like Bilby model the evenly sampled time-domain detector data $\bm{d}$ as a linear superposition of a possible GW signal $\bm{s}(\bm{\theta})$ and noise $\bm{n}$ (Ashton and others, 2019). Their goal is to identify a template $\bm{h}(\bm{\theta}\mid M)$ from a chosen waveform model $M$ (in our case, either TEOBResumS-Dalí or pTEOBResumS-Dalí) that best represents the signal, such that the residual $\bm{r}=\bm{d}-\bm{h}(\bm{\theta}\mid M)$ follows the noise’s statistical properties.
Generally, the noise $\bm{n}$ is assumed to be a zero-mean, wide-sense stationary Gaussian process with a covariance matrix $C_{ik}=\langle n_{i}n_{k}\rangle$. Here, $n_{i}$ and $n_{k}$ are the noise samples at discrete times $t_{i}$ and $t_{k}$, respectively. If we further assume that the data are periodic with a time period $T$, then the covariance matrix in the Fourier domain is diagonal,
$\tilde{C}_{jj}=\langle|\tilde{n}_{j}|^{2}\rangle=\frac{T}{2}S_{j},$ (2)
where $S_{j}$ is the one-sided noise power spectral density and
$\tilde{n}_{j}=\sum_{k=0}^{N-1}n_{k}\,e^{-2\pi ijk/N}$ (3)
is the discrete Fourier transform of $\bm{n}$ at frequency-bin index $j$, with $f_{s}=N/T$ the sampling frequency.
Under these assumptions, the single-detector log-likelihood ratio between the signal and noise assertion for the full analysis is (Finn, 1992; Veitch and Vecchio, 2010)
$\ln\mathcal{L}^{S}_{N}=-\frac{1}{2}\langle\bm{h}\mid\bm{h}\rangle+\langle\bm{d}\mid\bm{h}\rangle,$ (4)
where
$\langle\bm{a}\mid\bm{b}\rangle=\Re\sum_{j}\breve{a}_{j}^{\ast}\,\breve{b}_{j}$ (5)
is the noise-weighted inner product in the Fourier domain, and the whitened series is
$\breve{a}_{j}=\tilde{a}_{j}\sqrt{\frac{4}{TS_{j}}}.$ (6)
Equivalently, we may express the likelihood in the discrete time domain using the noise covariance matrix $C_{ik}$ (Isi and Farr, 2021). In our analysis, we compute $C$ from the autocorrelation function obtained by inverse Fourier transforming the PSD $S_{j}$. Assuming wide-sense stationarity, $C$ is Toeplitz and positive definite; we perform a Cholesky decomposition,
$C=L\,L^{\mathrm{T}},$ (7)
to define whitened time-domain quantities
$\bar{a}_{i}=\sum_{k}(L^{-1})_{ik}a_{k},$ (8)
Restricting the analysis to a time window $t_{\mathrm{start}},\,t_{\mathrm{end}}]$, we define the time-limited noise-weighted inner product
$\langle\bm{a}\mid\bm{b}\rangle_{[t_{\mathrm{start}},\,t_{\mathrm{end}}]}=\sum_{i}\bar{a}_{i}\,\bar{b}_{i},$ (9)
which we use to evaluate the log-likelihood ratio defined in Eq. (4) to probe the temporal structure of $(4,\pm 4)$ modes (Chandra and Calderón Bustillo, 2025).
Instead of performing a new stochastic sampling for this restricted window, we recompute the likelihood using Eq. (4) by resampling from the existing IMR posteriors (Rover et al., 2006; Ashton, 2025). This means the samples from the IMR analysis act as draws from a proposal distribution, while the restricted time-domain likelihood defines a new target distribution. The ratio of these two, expressed as a weight factor for each sample, reweights both the IMR posterior and the evidence, providing an approximation to the restricted-domain posterior without rerunning the full sampler.
This method is computationally efficient but comes with an important caveat: the true posterior from a restricted time-domain analysis is generally broader than that from the full IMR run, since the restricted likelihood omits part of the signal and weakens correlations among parameters. Its validity, therefore, depends on the proposal distribution adequately covering the high-probability region of the target distribution; any coverage gaps or sampling bias in the IMR analysis will be inherited by the resampled posterior. Given that our primary interest is the relative evidence, this procedure offers a fast, low-cost approximation. To ensure its robustness, we independently cross-checked the results by carrying out full analyses for $\Delta t^{\mathrm{geo}}_{\mathrm{start}}/M\in\{-5,\,0,\,5,\,10\}$, confirming consistency with the resampled estimates.
Figure 9: Left panel: Unfaithfulness $\bar{\mathcal{F}}$ between the NRSur7dq4 and TEOBResumS-Dalí waveform model for binary parameters consistent with the posteriors. The dashed and dotted lines show the mismatches obtained when no in-plane rotation ($\theta$) and azimuth ($\phi$) minimisations are performed. Different colors correspond to mismatches obtained considering models up to $\ell=2$ (green) and $\ell=4$ (pink). The vertical black line marks the critical threshold below which the models are indistinguishable, ensuring unbiased inference at the signal’s SNR. Right panel: Joint Posterior of mass and spin parameters for GW250114 inferred using the TEOBResumS-Dalí (red) and NRSur7dq4 (teal) waveform model. The posteriors broadly agree.
Comparison against surrogate models:
To assess the robustness of our GR baseline model, we compute mismatches between TEOBResumS-Dalí and NRSur7dq4 (Varma et al., 2019), a quasi-spherical multipolar model built by interpolating NR simulations and widely regarded as the most NR-faithful waveform currently available. We use the sky-maximized detection statistics of Harry et al. (2018), maximized over time and reference azimuth, as well as over an additional in-plane rotation angle Pratten and others (2021). We assume a lower frequency cutoff of $20$ Hz and an upper cutoff of $900$ Hz, employing the Advanced LIGO design sensitivity curve (Aasi and others, 2015). We find mismatches of $\bar{\mathcal{F}}\sim 10^{-4}-10^{-3}$ ($10^{-3}-10^{-2}$) between the two models when considering only the $\ell=2$ (including up to $\ell=4$) modes for systems drawn from the posterior of GW250114 (see left panel of Fig. 9). These numbers translate into critical SNR thresholds of $\rho_{\mathrm{crit}}\sim 70-100$ (203020-30) below which the two models should be indistinguishable Damour et al. (2011). Therefore, while the $\ell=2$ modes of the two models are mostly indistinguishable for GW250114, the inclusion of higher-order modes may in principle lead to measurable differences given the signal’s high SNR.
As a further validation, we repeated the parameter estimation (PE) using NRSur7dq4, and compared the results with those obtained from TEOBResumS-Dalí. The right panel of Figure 9 shows the posterior distributions of the mass and spin parameters, which are in broad agreement across the two models.
Toy Monte Carlo Proof of GW250114’s Isolated-Binary Merger Origin:
To demonstrate that GW250114 is less likely to have formed through a hierarchical merger process, we conduct a simple Monte Carlo study informed by the GWTC–3 population estimates (Abbott and others, 2023). We generate an initial BBH population by drawing component masses $m_{k}$ from the Power Law + Peak model, and spin magnitudes $\chi_{k}$ and tilt angles $\theta_{k}$ from the Default spin model (Talbot and Thrane, 2017, 2018). We label the BHs in this sample as “1G”, while noting that the observed GWTC–3 population may itself contain hierarchically assembled remnants; thus, this population does not represent a pure 1G population aka BHs formed directly via stellar-collapse.
We merge a fraction of this 1G+1G population to form second-generation (2G) BHs. Their final masses and spins are computed using NR-informed fits (Hofmann et al., 2016; Jiménez-Forteza et al., 2017), and their recoil speeds are obtained using fits from Gerosa and Kesden (2016). For subsequent steps, we retain only those remnants whose recoil speeds are below $500\,\mathrm{km\,s^{-1}}$, a threshold achievable only in a few host environments (e.g., AGN discs (Gerosa and Berti, 2017)); thus, our results provide optimistic upper bounds for hierarchical growth.
Next, we construct mixed-generation 1G+2G and 2G+2G binaries by pairing black holes from the 1G and 2G samples with equal probability. We draw their spin orientations from an isotropic distribution, as expected because dynamical pairing leads to random spin alignment. We subsequently merge these binaries to produce third-generation (3G) BHs and explore 1G+3G, 2G+3G, and 3G+3G pairings in a similar fashion.
For each population (1G+1G, 1G+2G, 2G+2G, …), we compute the probability using GW250114’s posteriors, obtained under the quasi-spherical BBHs interpretation, to belong to a given population. We find
$P_{\mathrm{1G+1G}}\sim 0.967,\quad P_{\mathrm{1G+2G}}\sim 0.032,\quad P_{\mathrm{2G+2G}}\sim 0.001.$
The much lower probabilities for hierarchical pairings arise because near-equal-mass, low-spin binaries ($q\to 1$, $\chi_{1,2}\lesssim 0.2$) tend to produce remnants with $\chi_{f}\gtrsim 0.5$, largely independent of spin orientation. Probabilities fall further for 1G+3G, 2G+3G, and 3G+3G mergers. We repeat this analysis multiple times and obtain similar numbers every time.
In summary, this simple experiment strongly favours a 1G+1G origin for GW250114. While idealised—neglecting detection probabilities, realistic retention fractions, pairing biases, and environment-specific dynamics—our assumptions already maximise the probability of hierarchical assembly.